OpenStudy (anonymous):

Let f(x) = sqrt(x^2 + 16) - 5 Find f(1+a).

TheSmartOne (thesmartone):

plug in 1+a in place of x

OpenStudy (anonymous):

yes i ended up with "a" which is wrong

OpenStudy (anonymous):

not sure what I did wrong

OpenStudy (anonymous):

\[f(x) = \sqrt{x^2 + 16} - 5\] \[f(1 + a) = \sqrt{(1 + a)^2 + 16} - 5 = ~a? ~~really?\] @PeterSheperd

OpenStudy (alexandervonhumboldt2):

\[f(1+a)=\sqrt{(1+a)^2+16}-5\]

TheSmartOne (thesmartone):

\[f(1+a)=\sqrt{(1+2a + a^2+16}-5\]

OpenStudy (alexandervonhumboldt2):

DIRECT ANSWER

TheSmartOne (thesmartone):

I haven't even completed it, lol

OpenStudy (anonymous):

ah i thought you could just cancel the exponent and take square root of 16

OpenStudy (anonymous):

Smart One ~__~ @TheSmartOne

OpenStudy (alexandervonhumboldt2):

not true lol

OpenStudy (alexandervonhumboldt2):

i would say that this cannot be simplified any further

OpenStudy (alexandervonhumboldt2):

all wut you can do is simplifying 1+16 and that is the end

TheSmartOne (thesmartone):

I just did one step, I didn't even do it all .-.

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @PeterShepherd ah i thought you could just cancel the exponent and take square root of 16 \(\color{blue}{\text{End of Quote}}\) nope you can't do that \[\rm \sqrt{(1+a)^2+16}-5 \cancel{=} \sqrt{(1+a)^2}+\sqrt{16} \] there is a plus sign between x^2 and 16 so it's not possible to splt them into two parts it's not like a multiplication